14. How much heat does it take to boil (100°C) 3.50 kg of water from room temperature (22°C)? (cwater = 4180 J/kg C)


1
Expert's answer
2021-04-29T11:01:09-0400

Solution.

"m=3.50kg;"

"t_1=22^oC;"

"t_2=100^oC;"

"c=4180J\/kgC;"

"Q-?;"

"Q=cm\Delta t;"

"Q=4180J/kgC\sdot3.50kg\sdot\sdot(100-22)^oC=14552J;"

Answer: "Q=14552J."